# Creating a Hybrid Rule and Neural Network Based Semantic Tagger using Silver Standard Data: the PyMUSAS framework for Multilingual Semantic Annotation

### v1

@@ Abstract
Word Sense Disambiguation (WSD) has been widely evaluated using the semantic frameworks of WordNet (Maru et al., 2022), BabelNet (Pasini et al., 2021), and the Oxford Dictionary of English (Gadetsky et al., 2018; Chang et al., 2018). However, for the UCREL Semantic Analysis System (USAS) framework, no open extensive evaluation has been performed beyond lexical coverage or single language evaluation. In this work, we perform the largest semantic tagging evaluation of the rule based system that uses the lexical resources in the USAS framework covering five different languages using four existing datasets and one novel Chinese dataset. We create a new silver labelled English dataset, to overcome the lack of manually tagged training data, that we train and evaluate various mono and multilingual neural models in both mono and cross-lingual evaluation setups with comparisons to their rule based counterparts, and show how a rule based system can be enhanced with a neural network model. The resulting neural network models, including the data they were trained on, the Chinese evaluation dataset, and all of the code have been released as open resources.
@@
Andrew Moorethanks: Corresponding authors. Johanna Vuorinen and the referenced Laura Löfberg refer to the same individual.∗1, Paul Rayson∗1, Dawn Archer2, Tim Czerniak3, Dawn Knight4,
@@
7Hubei University, China;
@@ 2.  USAS Tagset
The USAS tagset (Rayson et al., 2004; Löfberg and Rayson, 2019), which was originally derived from the Longman Lexicon of Contemporary English (McArthur, 1981), contains 21 major discourse fields that expand to 232 category labels. USAS has a hierarchical tagset that has up to three levels of sub-division. An example of the tagset is shown in Table 3. These 232 category labels can be enhanced with affixed symbols to indicate additional linguistic meaning, such as rarity (%@), gender (mf), antonyms (+-), and can be applied to both single words and Multi-Word Expressions (MWE). The semantic tags can be combined with double or triple membership through the use of a slash ("/"). For instance, in Table 2 the token "Vac" is tagged with "F2/O2", indicating the token is related to both Drinks (F2) and Objects (O2), which is correct as the text comes from a coffee related article [^8] Wikipedia article on the vacuum coffee maker (Vac pot): https://en.wikipedia.org/wiki/Vacuum_coffee_maker. Examples of semantically tagged tokens with the full USAS tagset can be seen in table 2, and a full explanation of the USAS tagset can be found in Archer et al. (2002).
@@ 3.2.  Evaluation Data
For the Welsh language, we used the entire dataset released in Ezeani et al. (2019) for evaluation. The dataset consists of 8 extracts from 4 diverse data sources; Kynulliad3131313https://kevindonnelly.org.uk/kynulliad3/ (Welsh Assembly proceedings), Meddalwedd [^14] https://techiaith.cymru/corpws/Moses/Meddalwedd/ (translations of software instructions), Kwici [^15] https://kevindonnelly.org.uk/kwici/ (Welsh Wikipedia articles), and LERBIML [^16] https://www.lancaster.ac.uk/fass/projects/biml/bimls3corpus.htm (multi-domain spoken corpora). In the original work with manually corrected POS tags, 90% of this dataset was used for training a joint POS and semantic tagger, while 10% was used for evaluation.
@@
The Chinese dataset that we have created is a manually tagged text from the “News Report” genre of the ToRCH2019 corpus (Jialei Li, Mingchen Sun, and Jiajin Xu, 2022), specifically about the 2019 military world games in Wuhan, China. This dataset was manually annotated following a three-stage procedure; 1. independent tagging, 2. independent reviews of their own tagging, and 3. reaching consensus between two trained researchers. By comparing the results of automatic and manual annotation, we propose recommendations for segmenting and annotating Chinese measure words, dates, and time expressions to enhance the precision of automated Chinese semantic annotation.
@@ 4.1.  Rule-Based Models
For each language we evaluate on, we start by using the existing resources and methods with PyMUSAS, i.e. each language has its own rule-based tagger along with the resources, as shown in Table 8202020All spaCy Transformer and large language models can be found at: https://spacy.io/models. The Irish lexicons come from Czerniak and Uí Dhonnchadha (2024) [^21] The Irish tagger from the original paper has been used here. A later improved version will be used in future testing., the English, Chinese and Finnish lexicons come from Piao et al. (2016), and the Welsh lexicons come from Piao et al. (2018) [^22] All the lexicons apart from Irish were acquired from: https://github.com/UCREL/Multilingual-USAS.
@@ 4.2.  Neural Models
The BEM model was chosen over more accurate and recent models from the WSD literature due to the model’s computational efficiency, while the more recent and accurate models (Barba et al., 2021; Zhang et al., 2022) are all cross-encoders. To use cross-encoders at inference time to disambiguate all words in a text would require running the PLM $n\times m$ times, whereby $n$ if the number of words in the given text to disambiguate and $m$ are the number of senses in the sense inventory compared to $m+1$ for the BEM model [^25] The BEM model can pre-compute the $m$ sense embeddings in this case disambiguating all words in a given text would only require running the context encoder PLM once.. The computational efficiency difference between bi-encoders and cross-encoders is highlighted in previous works (Reimers and Gurevych, 2019; Humeau et al., 2019).
@@ 6.  Results
For Chinese, it can be seen that, for both $n=1$ and $n=5$, the neural network models (NMulS and NMulB) outperformed both the rule-based and hybrid models due to the poor performance of the rule-based model. This is also the case for English when $n=5$.
@@ 7.  Conclusion and Future Work
Through the most extensive evaluation so far, with respect to number of languages, we have shown that either the hybrid or neural network based models across all languages outperform the existing rule-based methods, showing the advantages of these new neural network or hybrid based models for USAS tagging.
@@
In this work, we have released the first open manually annotated corpus for USAS semantic tagging in Chinese. We have also released the first silver labelled dataset in English that has been shown to be useful for training both mono and multilingual neural networks for USAS tagging. Future work could explore extending the silver labelled datasets into other languages using the same methodology of using existing rule-based USAS tagger, which will allow exploration of fine-tuning of language specific neural networks models.
@@ 10.  Bibliographical References
  • T. Chang, T. Chi, S. Tsai, and Y. (. Chen (2018) XSense: learning sense-separated sparse representations and textual definitions for explainable word sense networks. ArXiv abs/1809.03348. External Links: Link Cited by: Creating a Hybrid Rule and Neural Network Based Semantic Tagger using Silver Standard Data: the PyMUSAS framework for Multilingual Semantic Annotation.
@@
  • T. Czerniak and E. Uí Dhonnchadha (2024) Towards semantic tagging for Irish. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), N. Calzolari, M. Kan, V. Hoste, A. Lenci, S. Sakti, and N. Xue (Eds.), Torino, Italia, pp. 16643–16652. External Links: Link Cited by: §1, §3.2, §4.1, footnote 17, footnote 26.
@@
  • A. Gadetsky, I. Yakubovskiy, and D. Vetrov (2018) Conditional generators of words definitions. In Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), I. Gurevych and Y. Miyao (Eds.), Melbourne, Australia, pp. 266–271. External Links: Link, Document Cited by: Creating a Hybrid Rule and Neural Network Based Semantic Tagger using Silver Standard Data: the PyMUSAS framework for Multilingual Semantic Annotation.
@@
  • M. Maru, S. Conia, M. Bevilacqua, and R. Navigli (2022) Nibbling at the hard core of Word Sense Disambiguation. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), S. Muresan, P. Nakov, and A. Villavicencio (Eds.), Dublin, Ireland, pp. 4724–4737. External Links: Link, Document Cited by: Table 1, §1, Creating a Hybrid Rule and Neural Network Based Semantic Tagger using Silver Standard Data: the PyMUSAS framework for Multilingual Semantic Annotation.
@@
  • T. Pasini, A. Raganato, and R. Navigli (2021) XL-wsd: an extra-large and cross-lingual evaluation framework for word sense disambiguation. Proceedings of the AAAI Conference on Artificial Intelligence 35 (15), pp. 13648–13656. External Links: Link, Document Cited by: Table 1, §1, Creating a Hybrid Rule and Neural Network Based Semantic Tagger using Silver Standard Data: the PyMUSAS framework for Multilingual Semantic Annotation.

### v2

@@ Abstract
Word Sense Disambiguation (WSD) has been widely evaluated using the semantic frameworks of WordNet (Maru et al., 2022), BabelNet (Pasini et al., 2021), and the Oxford Dictionary of English (gadetsky-etal-2018-conditional; Chang2018xSenseLS). However, for the UCREL Semantic Analysis System (USAS) framework, no open extensive evaluation has been performed beyond lexical coverage or single language evaluation. In this work, we perform the largest semantic tagging evaluation of the rule based system that uses the lexical resources in the USAS framework covering five different languages using four existing datasets and one novel Chinese dataset. We create a new silver labelled English dataset, to overcome the lack of manually tagged training data, that we train and evaluate various mono and multilingual neural models in both mono and cross-lingual evaluation setups with comparisons to their rule based counterparts, and show how a rule based system can be enhanced with a neural network model. The resulting neural network models, including the data they were trained on, the Chinese evaluation dataset, and all of the code have been released as open resources.
@@
Andrew Moorethanks: Corresponding authors. Johanna Vuorinen and the referenced Laura Löfberg refer to the same individual.∗1, Paul Rayson∗1,8, Dawn Archer2, Tim Czerniak3, Dawn Knight4,
@@
7Hubei University, China; 8VinUniversity, Vietnam
@@ 2.  USAS Tagset
The USAS tagset (Rayson et al., 2004; Löfberg and Rayson, 2019), which was originally derived from the Longman Lexicon of Contemporary English (McArthur, 1981), contains 21 major discourse fields that expand to 232 category labels. USAS has a hierarchical tagset that has up to three levels of sub-division. An example of the tagset is shown in Table 3. These 232 category labels can be enhanced with affixed symbols to indicate additional linguistic meaning, such as rarity (%@), gender (mf), antonyms (+-), and can be applied to both single words and Multi-Word Expressions (MWE). The semantic tags can be combined with double or triple membership through the use of a slash ("/"). For instance, in Table 2 the token "Coffee" is tagged with "F2/O2", indicating the token is related to both Drinks (F2) and Objects (O2), which is correct. Examples of semantically tagged tokens with the full USAS tagset can be seen in table 2, and a full explanation of the USAS tagset can be found in Archer et al. (2002).
@@
O1.2 Substances and materials generally: Liquid
O1.3 Substances and materials generally: Gas
@@ 3.2.  Evaluation Data
For the Welsh language, we used the entire dataset released in Ezeani et al. (2019) for evaluation. The dataset consists of 8 extracts from 4 diverse data sources; Kynulliad3121212https://kevindonnelly.org.uk/kynulliad3/ (Welsh Assembly proceedings), Meddalwedd [^13] https://techiaith.cymru/corpws/Moses/Meddalwedd/ (translations of software instructions), Kwici [^14] https://kevindonnelly.org.uk/kwici/ (Welsh Wikipedia articles), and LERBIML [^15] https://www.lancaster.ac.uk/fass/projects/biml/bimls3corpus.htm (multi-domain spoken corpora). In the original work with manually corrected POS tags, 90% of this dataset was used for training a joint POS and semantic tagger, while 10% was used for evaluation.
@@
The Chinese dataset that we have created is a manually tagged text from the “News Report” genre of the ToRCH2019 corpus (Jialei Li, Mingchen Sun, and Jiajin Xu, 2022), specifically about the 2019 military world games in Wuhan, China. This dataset was manually annotated following a three-stage procedure; 1. independent tagging, 2. independent reviews of their own tagging, and 3. reaching consensus between two trained researchers.
@@ 4.1.  Rule-Based Models
For each language we evaluate on, we start by using the existing resources and methods with PyMUSAS, i.e. each language has its own rule-based tagger along with the resources, as shown in Table 8191919All spaCy Transformer and large language models can be found at: https://spacy.io/models. The Irish lexicons come from Czerniak and Uí Dhonnchadha (2024) [^20] The Irish tagger from the original paper has been used here. A later improved version will be used in future testing., the English, Chinese and Finnish lexicons come from Piao et al. (2016), and the Welsh lexicons come from Piao et al. (2018) [^21] All the lexicons apart from Irish are available from: https://github.com/UCREL/Multilingual-USAS.
@@ 4.2.  Neural Models
The BEM model was chosen over more accurate and recent models from the WSD literature due to the model’s computational efficiency, while the more recent and accurate models (Barba et al., 2021; Zhang et al., 2022) are all cross-encoders. To use cross-encoders at inference time to disambiguate all words in a text would require running the PLM $n\times m$ times, whereby $n$ is the number of words in the given text to disambiguate and $m$ are the number of senses in the sense inventory compared to $m+1$ for the BEM model [^24] The BEM model can pre-compute the $m$ sense embeddings in this case disambiguating all words in a given text would only require running the context encoder PLM once.. The computational efficiency difference between bi-encoders and cross-encoders is highlighted in previous works (Reimers and Gurevych, 2019; Humeau et al., 2019).
@@ 6.  Results
For Chinese, it can be seen that, for both $n=1$ and $n=5$, the neural network models (NMulS and NMulB) outperformed both the rule-based and hybrid models due to the poor performance of the rule-based model. This is also the case for English when $n=5$. This highlights a drawback of the hybrid model whereby if the recall of the rule-based model is high and the precision is low then the hybrid model will be limited by the rule-based model’s performance. It will be less likely to fall back to the neural network model due to the rule-based model’s high recall. This, we hypothesise is the reason why there is little difference in performance between the rule-based model and the hybrid model for English for both values of $n$.
@@ 7.  Conclusion and Future Work
Through the most extensive evaluation so far, with respect to number of languages, we have shown that either the hybrid or neural network based models across all languages outperform the existing rule-based methods, showing the advantages of these new neural network or hybrid based models for USAS tagging. In future work, we will evaluate the hybrid model further to disentangle the performance gains and losses from the neural network model and rule-based model respectively. We will also evaluate the neural network and rule-based models on tokens that are defined in the rule-based model’s lexicon(s) so that we can better isolate the disambiguate performance of the models, rather than evaluating both coverage and disambiguate, in addition this evaluation will inform us of whether neural network models can go beyond the performance of the rule-based model that it learnt from.
@@
In this work, we have released the first open manually annotated corpus for USAS semantic tagging in Chinese. Our initial qualitative comparison of the rule-based Chinese model with the manually annotated Chinese corpus revealed that Chinese measure words, dates, and time expressions are semantic categories where the model can be improved. These categories will be monitored and systematically evaluated in future work.
We have also released the first silver labelled dataset in English that has been shown to be useful for training both mono and multilingual neural networks for USAS tagging. Future work could explore extending the silver labelled datasets into other languages using the same methodology of using existing rule-based USAS tagger, which will allow exploration of fine-tuning of language specific neural networks models.
@@ 9.  Acknowledgements
We are grateful to the UCREL research group at Lancaster University for providing us access to the Hex computer cluster Vidler and Rayson (2024) where the experiments were processed.
@@ 10.  Bibliographical References
  • T. Czerniak and E. Uí Dhonnchadha (2024) Towards semantic tagging for Irish. In Proceedings of the 2024 Joint International Conference on Computational Linguistics, Language Resources and Evaluation (LREC-COLING 2024), N. Calzolari, M. Kan, V. Hoste, A. Lenci, S. Sakti, and N. Xue (Eds.), Torino, Italia, pp. 16643–16652. External Links: Link Cited by: §1, §3.2, §4.1, footnote 16, footnote 25.
@@
  • M. Maru, S. Conia, M. Bevilacqua, and R. Navigli (2022) Nibbling at the hard core of Word Sense Disambiguation. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), S. Muresan, P. Nakov, and A. Villavicencio (Eds.), Dublin, Ireland, pp. 4724–4737. External Links: Link, Document Cited by: Table 1, §1.
@@
  • T. Pasini, A. Raganato, and R. Navigli (2021) XL-wsd: an extra-large and cross-lingual evaluation framework for word sense disambiguation. Proceedings of the AAAI Conference on Artificial Intelligence 35 (15), pp. 13648–13656. External Links: Link, Document Cited by: Table 1, §1.
@@
  • J. Vidler and P. Rayson (2024) UCREL - Hex; a shared, hybrid multiprocessor system. Note: https://github.com/UCREL/hexAccessed: 2024 Cited by: §9.